OpenStudy (anonymous):

slope of this line -9x+12y=-9

OpenStudy (mertsj):

Solve the equation for y

OpenStudy (mertsj):

The coefficient of x will then be the slope.

OpenStudy (shamil98):

put the equation into slope-intercept form \[y=mx + b\] where m is your slope

OpenStudy (shamil98):

-9x + 12y = 9 12y = 9x + 9 divide everything by 12 y = 9/12 x + 9/12 you can simplify the fractions so do that. y = 3/4 x + 3/4

OpenStudy (anonymous):

Thanks..so .75 is the slope?

OpenStudy (anonymous):

@rachelparent you can just leave as 3/4

OpenStudy (anonymous):

ok thank you